OpenStudy (anonymous):

Use LaGrange multipliers to solve the following problem: A closed box is to be constructed with volume 750in^3. The material for the top and bottom costs 3 cents per square inch, the material for the front and back costs 6 cents per square inch, and the material for two ends costs 9 cents per square inch. What dimensions should the box have to minimize the cost of the construction materials? What is the minimum cost per box?

OpenStudy (anonymous):

 Here are your equations and constraints. \[V=xyz=750\] \[A=2xy+2xz+2yz\] \[C=3(2xy)+6(2xz)+9(2yz)\]

OpenStudy (anonymous):

The cost \(C\) is minimized when the amount of material \(A\) is minimized. Lagrange multipliers gives \[\begin{cases}\nabla A(x,y,z)=\lambda\nabla V(x,y,z)\\ V(x,y,z)=750\end{cases}\] or \[\begin{cases} 2y+2z=\lambda yz\\ 2x+2z=\lambda xz\\ 2x+2y=\lambda xy\\ xyz=750 \end{cases}\] You need to find \(x,y,z\) that meet these conditions.

OpenStudy (anonymous):

Ok so now just solve the system of equations and plug the answers into the original equation to find the minimum?

OpenStudy (anonymous):

Yes. You can assume \(\lambda\not=0\).